In what proportion should 95% alcohol be mixed with 30% alcohol to make 70% alcohol?
Answer by Earlsdon(6294)

Try this:
Let x = the amount of 95% alcohol mixture required.
Let y = the amount of 30% alcohol mixture required.
We can write:
0.95x+0.3y = 0.7(x+y) Expand the right side.
o.95x+0.3y = 0.7x+0.7y Combine like-terms.
0.25x = 0.4y
Since we want the ratio of the amount of 95% alcohol mixture (x) to the amount of 35% alcohol mixture (y), which is x/y, we can write:
x/y = 0.4/0.25 Simplifying this we get:
x/y = 8/5
So the proportion is 8 units of 95% alcohol mixture to 5 units of 30% alcohol mixture to make 8+5 = 13 units of 70% alcohol mixture.
Check:
8(0.95)+5(0.3) = 13(0.7)
7.6+1.5 = 9.1
9.1 = 9.1